# PhyWorldBench: A Comprehensive Evaluation of Physical Realism in Text-to-Video Models

### v1

@@ Abstract
Machine Learning, ICML
[…]
figure Overview of  PhyWorldBench. The benchmark follows a structured design, starting with 10 main physics categories, derived from physics literature and expert consultations. Each category is divided into 5 subcategories, capturing different aspects. Under each subcategory, 7 scenarios are created, with 3 prompt variations per scenario to provide varying levels of detail and complexity. The figure presents the benchmark structure, showcasing the 10 main categories and their corresponding 5 subcategories.
@@ 1 Introduction
The field of video generation has made remarkable progress, with models producing visually compelling and often photorealistic outputs. These advances have enabled transformative applications across industries such as entertainment, education, and scientific visualization. However, despite their visual fidelity, do video generation models truly understand the laws of physics in the real world? To answer this question, we introduce PhyWorldBench, a rigorous benchmark designed to evaluate how well video generation models can simulate real-world physics. As illustrated in PhyWorldBench: A Comprehensive Evaluation of Physical Realism in Text-to-Video Models, PhyWorldBench systematically tests models across multiple levels of physical phenomena, from fundamental concepts like object motion to complex dynamics, including rigid body interactions and human/animal motion. Additionally, we propose a novel Anti-Physics category, where prompts deliberately violate real-world physics. On one hand, this design verifies whether models genuinely understand physical laws—rather than merely reproducing patterns from real-world training data. On the other hand, anti-physics content itself holds practical value in creative applications, where imaginative or otherwise impossible scenarios are beneficial.
@@
To examine the current status of video generation models and provide a detailed analysis, we selected five proprietary models—Sora-Turbo (OpenAI, 2024), Gen-3 (Runway Team, 2024), Kling 1.6 (KlingAI, 2024), Pika 2.0 (Pika Labs Team, 2024), and Luma (Luma AI Team, 2024)—along with seven open-source models: Hunyuan 720p (Kong et al., 2024), Open-Sora 2.0 (Peng et al., 2025), Open-Sora-Plan 1.3 (Lin et al., 2024), CogVideo 14b (Hong et al., 2022), Step-video-T2V (Ma et al., 2025), Wanx-2.1 (WanTeam et al., 2025), and LTX-Video (HaCohen et al., 2024). We generated 12,600 videos to evaluate SOTA models and analyze their ability to simulate real-world physics. We identify challenges, difficult scenarios, and key physics categories while proposing a structured approach to improve physical realism through prompt design.
@@ 2 Related Work
Benchmarks for Text-to-Video Generation. Proprietary video generators (OpenAI, 2024; KlingAI, 2024; Pika Labs Team, 2024; Runway Team, 2024; HailuoAI, 2024; Luma AI Team, 2024) achieve striking quality and temporal coherence but remain opaque. Open-source counterparts (Peng et al., 2025; Lin et al., 2024; Hong et al., 2022; Wang et al., 2023; HaCohen et al., 2024; Kong et al., 2024; Agarwal et al., 2025) enable reproducible research and flexible benchmarking. The development of Text-to-Video (T2V) Generation models has been accelerated by benchmarks that evaluate performance across diverse aspects like quality, realism, and compositionality. VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024) provide comprehensive frameworks for assessing video generation models, focusing on metrics such as diversity, temporal coherence, and scalability. Besides, T2V-CompBench (Sun et al., 2024) emphasizes compositionality, ensuring generated videos align with intricate textual prompts, a key challenge in real-world applications. Recent works emphasize the importance of physical plausibility in video generation. VideoPhy (Bansal et al., 2024) and PhyGenBench (Meng et al., 2024) assess models’ adherence to physical commonsense, while DEVIL (Liao et al., 2024) evaluates dynamic realism, including motion and temporal consistency. Physics-IQ (Motameda et al., 2025) further tests understanding of fundamental principles like fluid dynamics and thermodynamics, highlighting the demand for physics-aware video generation.
@@ 3.1 Dataset Creation Pipeline
We identify broad categories—such as kinematics, rigid body dynamics, fluid behavior, optics, and thermal dynamics—as depicted in the main physics categories in PhyWorldBench: A Comprehensive Evaluation of Physical Realism in Text-to-Video Models. Each category is systematically divided into five subcategories to ensure comprehensive coverage of physics principles from multiple perspectives.
@@
We first use state-of-the-art large language models (LLMs), specifically GPT-4o (OpenAI Team, 2024) and Gemini-1.5-Pro (Gemini Team, 2024) to generate 20 Event Prompts under the predefined category and subcategory. Then we ask humans to select 7 out of them based on diversity and physics representation. Then human experts meticulously review and validate the generated prompts, providing feedback to iteratively enhance their quality and relevance to the corresponding physics subcategory, ensuring a balanced and representative assessment across various aspects of physical understanding. Furthermore, an ethical review is conducted to ensure that all prompts are fair, unbiased, and free from potentially harmful content. Based on the Event Prompt, human annotators create a Physics-Enhanced Prompt by enriching it with physical consequences. Detailed Narrative Prompt is generated by LLM based on Event Prompt using query in Table 11. This rigorous verification process results in a refined set of 1,050 high-quality prompts, systematically organized into main and subcategories.
@@
Statistic VideoPhy PhyGenBench Physics-IQ PhyWorldBench
Physics Categories 5 27 5 50
Prompts 688 160 396 1050
[…]
@@ 3.2 Statistics of the dataset
We compare our benchmark with existing physics-focused T2V benchmarks (Bansal et al., 2024; Meng et al., 2024), with a detailed statistical comparison presented in Table 1. PhyWorldBench offers the most extensive coverage in terms of the number of physics categories, curated prompts, human annotations, and evaluated T2V videos.
@@ 3.3 MLLM as Zero Shot Physics Evaluator for Generated Videos
Human evaluation, while flexible, is costly at scale. Existing metrics often lack generalizability and struggle with the diversity in PhyWorldBench. (See Appendix C.) MLLMs like GPT-o1 (OpenAI Team, 2024) and Gemini (Gemini Team, 2024) show promise for video evaluation but fall short in reliably assessing physical realism (Bansal et al., 2024; He et al., 2024; Meng et al., 2024). We find that these models tend to rationalize unrealistic content, but their accuracy improves when informed that the video is AI-generated, as illustrated in Figure 5. We hypothesize that this improvement stems from MLLMs being primarily trained on real-world videos, leading them to inherently justify observed phenomena unless prompted otherwise. Inspired by this, we propose a simple yet effective method, Context-Aware Prompt (CAP), which explicitly informs the MLLM that the video is generated rather than real. This approach enables the MLLM to provide a more accurate assessment of the video’s physical realism. Critically, CAP does not assume videos are incorrect by default but instead uses a structured chain‑of‑thought prompt—first asking the model to describe the video and then to reason through any potential physics issues—and only labels a video as incorrect if it clearly violates the category‑specific standards. We validate CAP neutrality by measuring false positive rates in physics correct videos and observe only a 0.001 absolute change in semantic adhesion (0.188 to 0.187) and a 0.014 change in physical commonsense (0.172 to 0.186), showing that it does not significantly increase misclassification of correct videos.
Besides, inspired by Chain-of-Thought (Wei et al., 2022), we design a two-step prompting strategy, where we first ask the MLLM to give a thorough description of the video including Object, Event, and Observations, then continue to prompt the MLLM for the final “Yes/No” answer for video quality. Please refer to Appendix I for prompts used in the two steps. We found this leads to consistent improvement in various MLLMs. In Table 2, we present the ROC AUC calculations comparing CAP with human results on PhyWorldBench with videos generated by models in Figure 1, where a random guess would achieve a score of 50. We evenly sample 8 frames. In this evaluation, CAP is powered by GPT-o1. CAP significantly outperforms other methods, particularly in Physics Commonsense, where it achieves an absolute performance boost of 13.5 compared to the standard GPT-o1. We find CAP also achieves great performance when equipping other proprietary or open-sourced MLLM. Table 2 shows the ablation study of various components in CAP. “without CoT” means we directly ask the MLLM to give the answer of the question, and “without Context” means we do not inform MLLM that the video could be of low quality. Please refer Appendix C for performance of our method on other MLLM and analysis.
@@ 4 Video Generation Results and Analysis
We evaluate both closed-source models including Sora-Turbo (OpenAI, 2024), Gen-3 (Runway Team, 2024), Kling 1.6 (KlingAI, 2024), Pika 2.0 (Pika Labs Team, 2024), Luma (Luma AI Team, 2024), and open-source models including Hunyuan 720p (Kong et al., 2024), Open-Sora 2.0 (Peng et al., 2025), Open-Sora-Plan 1.3 (Lin et al., 2024), CogVideo 14b (Hong et al., 2022), LTX-Video (HaCohen et al., 2024), Step-video-T2V (Ma et al., 2025), Wanx-2.1 (WanTeam et al., 2025). Please check Appendix D for the implementation details. We generate 1050 videos for each model. We conduct a human evaluation of generated videos using Amazon Mechanical Turk. We first introduce model performance and model-wise analysis and then break down the challenge for the video generation model to achieve real-world physics.
@@ 4.1 Model Performance
Table 3 presents a comprehensive evaluation of twelve video generation models. We report the percentage of videos that satisfy both semantic adherence and physical commonsense criteria (Both). The Overall Performance column provides an aggregate score for each model, averaged across 10 physics categories. Pika 2.0 achieves the best overall performance across all models. In open-sourced models, Hunyuan 720p achieves the best performance on PC, Wanx-2.1 achieves the best performance on Both, and CogVideo achieves the best SA performance. It is worth noting that they even achieve a better performance compared with some of the proprietary. Please check  Appendix L for Leaderboard.
@@ 4.2 Model Specific Analysis
Our evaluation reveals a trade-off between realism, stylization, and consistency in text-to-video models. Wanx, Step-Video-T2V, Gen-3, Hunyuan, Kling, and Sora produce the most realistic videos, while Pika and Luma add stylized lighting and colors. Open-Sora and LTX-Video struggle with realism, often generating low-fidelity or distorted outputs. Object stability varies—CogVideo, Open-Sora, and LTX-Video frequently produce warped shapes, while Gen-3, Hunyuan, and Kling maintain more stable forms. Motion realism also differs: Kling and Gen-3 generate smooth motion, Open-Sora-Plan and Luma introduce slow-motion effects, and CogVideo and Open-Sora exhibit jittery, erratic movement. For a detailed analysis, see Appendix J.
@@ 4.3 Effect of Physics and Prompt Enhancement
Understanding real-world physics is essential for text-to-video models. Event-based prompts describe actions, while physics-enhanced prompts include natural consequences. A model’s ability to generate realistic videos based on these prompts reflects its real-world understanding. As proprietary models conceal their generation processes, we cannot verify prompt refinement. Instead, we analyze prompt types using open-source models, where prompt integrity can be manually ensured. In Table 4, we compare model performance across different prompt types to gauge the effects of physics-aware prompting and refinement. While refinement improves clarity, object detail, and aesthetics, it does not notably enhance physics accuracy. The persistent inconsistencies suggest that these refinements primarily address surface-level quality, leaving deeper physics limitations unresolved. On the other hand, physics-enhanced prompts often yield better results, reflecting the models’ strong prompt-following ability—even though they still struggle with true physical comprehension. Consequently, we propose a straightforward receipt: explicitly integrate physical phenomena into the prompt to nudge the model toward more realistic outcomes.
@@ 4.4 Anti-Physics Analysis
As shown in Table 3, all models show a performance decline from Fundamental to Composite to Anti-Physics categories, reflecting the challenge of generating complex physical phenomena. While they capture basic physics, simulating intricate interactions and “unphysical” scenes remains difficult, highlighting a gap in physics adherence. Our analysis found that failures often stem from models interpreting prompts in a more “reasonable” way rather than attempting and failing. For example, when prompted with “A glass sits on the table, and the wine in the glass has reversed gravity,” the model generates a still image, adhering to real-world physics instead of the prompt. This suggests that models prioritize mimicking training data over true physics understanding.
@@ 4.5 Failure to Handle Erratic Visual Changes
Certain physical events inherently result in erratic visual changes, such as collisions, breaking, or sudden state transitions. However, current video generation models struggle to accurately depict these scenarios. In the third example of Figure 9, a glass cup falls from a table, an event that should cause it to shatter upon impact, leading to a sudden and complex visual transformation. Instead, the model rationalizes the action, generating a video where the glass remains intact and in the same vertical position after falling, completely ignoring the expected breakage and fragmentation. This highlights a fundamental limitation: when an event is supposed to cause an erratic visual change, current AI models fail to reflect the correct physical outcome. Rather than simulating realistic disruptions in structure, motion, or object state, the models tend to default to continuous, smoothed animations, avoiding abrupt transformations that are critical for representing real-world physics.

### v2

@@ Abstract
Figure 1: Overview of  PhyWorldBench. The benchmark follows a structured design, starting with 10 main physics categories, derived from physics literature and expert consultations. Each category is divided into 5 subcategories, capturing different aspects. Under each subcategory, 7 scenarios are created, with 3 prompt variations per scenario to provide varying levels of detail and complexity. The figure presents the benchmark structure, showcasing the 10 main categories and their corresponding 5 subcategories.
@@ 1 Introduction
The field of video generation has made remarkable progress, with models producing visually compelling and often photorealistic outputs. These advances have enabled transformative applications across industries such as entertainment, education, and scientific visualization. However, despite their visual fidelity, do video generation models truly understand the laws of physics in the real world? To answer this question, we introduce PhyWorldBench, a rigorous benchmark designed to evaluate how well video generation models can simulate real-world physics. As illustrated in Figure 1, PhyWorldBench systematically tests models across multiple levels of physical phenomena, from fundamental concepts like object motion to complex dynamics, including rigid body interactions and human/animal motion. Additionally, we propose a novel Anti-Physics category, where prompts deliberately violate real-world physics. On one hand, this design verifies whether models genuinely understand physical laws—rather than merely reproducing patterns from real-world training data. On the other hand, anti-physics content itself holds practical value in creative applications, where imaginative or otherwise impossible scenarios are beneficial.
@@
To examine the current status of video generation models and provide a detailed analysis, we selected five proprietary models—Sora-Turbo (OpenAI, 2024), Gen-3 (Runway Team, 2024), Kling 1.6 (KlingAI, 2024), Pika 2.0 (Pika Labs Team, 2024), and Luma (Luma AI Team, 2024)—along with seven open-source models: Hunyuan 720p (Kong et al., 2024), Open-Sora 2.0 (Peng et al., 2025), Open-Sora-Plan 1.3 (Lin et al., 2024), CogVideoX-1.5 (Hong et al., 2022), Step-video-T2V (Ma et al., 2025), Wanx-2.1 (WanTeam et al., 2025), and LTX-Video (HaCohen et al., 2024). We generated 12,600 videos to evaluate SOTA models and analyze their ability to simulate real-world physics. We identify challenges, difficult scenarios, and key physics categories while proposing a structured approach to improve physical realism through prompt design.
@@ 2 Related Work
Benchmarks for Text-to-Video Generation. Proprietary video generators (OpenAI, 2024; KlingAI, 2024; Pika Labs Team, 2024; Runway Team, 2024; HailuoAI, 2024; Luma AI Team, 2024) achieve striking quality and temporal coherence but remain opaque. Open-source counterparts (Peng et al., 2025; Lin et al., 2024; Hong et al., 2022; Wang et al., 2023; HaCohen et al., 2024; Kong et al., 2024; Agarwal et al., 2025) enable reproducible research and flexible benchmarking. The development of Text-to-Video (T2V) Generation models has been accelerated by benchmarks that evaluate performance across diverse aspects like quality, realism, and compositionality. VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024) provide comprehensive frameworks for assessing video generation models, focusing on metrics such as diversity, temporal coherence, and scalability. Besides, T2V-CompBench (Sun et al., 2024) emphasizes compositionality, ensuring generated videos align with intricate textual prompts, a key challenge in real-world applications. Recent works emphasize the importance of physical plausibility in video generation. Morpheus Zhang et al. (2025a) focuses on experimental physics. VideoPhy (Bansal et al., 2024) and PhyGenBench (Meng et al., 2024) assess models’ adherence to physical commonsense, while DEVIL (Liao et al., 2024) and VideoPhy-2 (Bansal et al., 2025) evaluates dynamic realism, including motion and temporal consistency. Physics-IQ (Motameda et al., 2025) further tests understanding of fundamental principles like fluid dynamics and thermodynamics, highlighting the demand for physics-aware video generation. Many physics-focused video generation models are also built (Zhang et al., 2025b; Yuan et al., 2025; Wang et al., 2025).
@@ 3.1 Dataset Creation Pipeline
We identify broad categories—such as kinematics, rigid body dynamics, fluid behavior, optics, and thermal dynamics—as depicted in the main physics categories in Figure 1. Each category is systematically divided into five subcategories to ensure comprehensive coverage of physics principles from multiple perspectives.
@@
We first use state-of-the-art large language models (LLMs), specifically GPT-4o (OpenAI Team, 2024) and Gemini-1.5-Pro (Gemini Team, 2024) to generate 20 Event Prompts under the predefined category and subcategory. Then we ask humans to select 7 out of them based on diversity and physics representation. Then human experts meticulously review and validate the generated prompts, providing feedback to iteratively enhance their quality and relevance to the corresponding physics subcategory, ensuring a balanced and representative assessment across various aspects of physical understanding. Furthermore, an ethical review is conducted to ensure that all prompts are fair, unbiased, and free from potentially harmful content. Based on the Event Prompt, human annotators create a Physics-Enhanced Prompt by enriching it with physical consequences. Detailed Narrative Prompt is generated by LLM based on Event Prompt using query in Tab. 12. This rigorous verification process results in a refined set of 1,050 high-quality prompts, systematically organized into main and subcategories.
@@
Statistic VideoPhy PhyGenBench Physics-IQ T2VPhysBench PhyWorldBench
Physics Categories 5 27 5 3 50
Prompts 688 160 396 84 1050
Step 3: Standard Creation. We adopt a Yes/No evaluation metric to assess whether a model can generate videos that accurately align with a given prompt. This metric is chosen for its simplicity and effectiveness and for allowing for an objective assessment while minimizing subjectivity in evaluation. As shown in Figure 5, we define two types of evaluation standards: Basic Standards and Key Standards. The Basic Standards specify the essential object(s) that should appear in the video and ensure that the main action or event is depicted. Meanwhile, the Key Standards describe the key physical phenomena that would occur if the event took place in the real world. The annotators are instructed to focus on the natural physical consequences of the action and to be explicit and concise in their annotations. By following these guidelines, we ensure that the evaluation standards remain clear and objective. Following  (Bansal et al., 2024; Meng et al., 2024), we use Semantic Adherence (SA) and Physical Commonsense (PC) to evaluate video performance. SA checks if both “objects” and “Action/Event” align with video, while PC assesses adherence to real-world physics and it checks when all “Key Standards” are satisfied. Annotated as $\text{SA},\text{PC}\in{0,1}$ (1 indicating proper grounding), we report the fraction of videos satisfying a success when $\text{SA}=1$, $\text{PC}=1$, and both jointly.
@@ 3.2 Statistics of the dataset
We compare our benchmark with existing physics-focused T2V benchmarks (Bansal et al., 2024; Meng et al., 2024; Guo et al., 2025; Motameda et al., 2025), with a detailed statistical comparison presented in Tab. 1. PhyWorldBench offers the most extensive coverage in terms of the number of physics categories, curated prompts, human annotations, and evaluated T2V videos.
@@ 3.3 MLLM as Zero Shot Physics Evaluator for Generated Videos
Human evaluation, while flexible, is costly at scale. Existing metrics often lack generalizability and struggle with the diversity in PhyWorldBench. (See Appx. D.) MLLMs like GPT-o1 (OpenAI Team, 2024) and Gemini (Gemini Team, 2024) show promise for video evaluation but fall short in reliably assessing physical realism (Bansal et al., 2024; He et al., 2024; Meng et al., 2024). We find that these models tend to rationalize unrealistic content, but their accuracy improves when informed that the video is AI-generated, as illustrated in Figure 6. We hypothesize that this improvement stems from MLLMs being primarily trained on real-world videos, leading them to inherently justify observed phenomena unless prompted otherwise. Inspired by this, we propose a simple yet effective method, Context-Aware Prompt (CAP), which explicitly informs the MLLM that the video is generated rather than real. This approach enables the MLLM to provide a more accurate assessment of the video’s physical realism. Critically, CAP does not assume videos are incorrect by default but instead uses a structured chain‑of‑thought prompt—first asking the model to describe the video and then to reason through any potential physics issues—and only labels a video as incorrect if it clearly violates the category‑specific standards. We validate CAP neutrality by measuring false positive rates in physics correct videos and observe only a 0.001 absolute change in semantic adhesion (0.188 to 0.187) and a 0.014 change in physical commonsense (0.172 to 0.186), showing that it does not significantly increase misclassification of correct videos.
Besides, inspired by Chain-of-Thought (Wei et al., 2022), we design a two-step prompting strategy, where we first ask the MLLM to give a thorough description of the video including Object, Event, and Observations, then continue to prompt the MLLM for the final “Yes/No” answer for video quality. Please refer to Appx. K for prompts used in the two steps. We found this leads to consistent improvement in various MLLMs. In Tab. 5, we present the ROC AUC calculations comparing CAP with human results on PhyWorldBench with videos generated by models in Figure 2, where a random guess would achieve a score of 50. We evenly sample 8 frames. In this evaluation, CAP is powered by GPT-o1. CAP significantly outperforms other methods, particularly in Physics Commonsense, where it achieves an absolute performance boost of 13.5 compared to the standard GPT-o1. We find CAP also achieves great performance when equipping other proprietary or open-sourced MLLM. The bottom of Tab. 5 shows the ablation study of various components in CAP. “without CoT” means we directly ask the MLLM to give the answer of the question, and “without Context” means we do not inform MLLM that the video could be of low quality. Please refer Appx. D for performance of our method on other MLLM and analysis.
@@ 4 Video Generation Results and Analysis
We evaluate both closed-source models including Sora-Turbo (OpenAI, 2024), Gen-3 (Runway Team, 2024), Kling 1.6 (KlingAI, 2024), Pika 2.0 (Pika Labs Team, 2024), Luma (Luma AI Team, 2024), and open-source models including Hunyuan 720p (Kong et al., 2024), Open-Sora 2.0 (Peng et al., 2025), Open-Sora-Plan 1.3 (Lin et al., 2024), CogVideoX-1.5 (Hong et al., 2022), LTX-Video (HaCohen et al., 2024), Step-video-T2V (Ma et al., 2025), Wanx-2.1 (WanTeam et al., 2025). Please check Appx. E for the implementation details. We generate 1050 videos for each model. We conduct a human evaluation of generated videos using Amazon Mechanical Turk. We first introduce model performance and model-wise analysis and then break down the challenge for the video generation model to achieve real-world physics.
@@ 4.1 Model Performance
Tab. 3 presents a comprehensive evaluation of twelve video generation models. We report the percentage of videos that satisfy both semantic adherence and physical commonsense criteria (Both). The Overall Performance column provides an aggregate score for each model, averaged across 10 physics categories. Pika 2.0 achieves the best overall performance across all models. In open-sourced models, Hunyuan 720p achieves the best performance on PC, Wanx-2.1 achieves the best performance on Both, and CogVideoX-1.5 achieves the best SA performance. It is worth noting that they even achieve a better performance compared with some of the proprietary. Please check Appx. P for Leaderboard.
@@ 4.2 Model Specific Analysis
Our evaluation reveals a trade-off between realism, stylization, and consistency in text-to-video models. Wanx, Step-Video-T2V, Gen-3, Hunyuan, Kling, and Sora produce the most realistic videos, while Pika and Luma add stylized lighting and colors. Open-Sora and LTX-Video struggle with realism, often generating low-fidelity or distorted outputs. Object stability varies—CogVideoX-1.5, Open-Sora, and LTX-Video frequently produce warped shapes, while Gen-3, Hunyuan, and Kling maintain more stable forms. Motion realism also differs: Kling and Gen-3 generate smooth motion, Open-Sora-Plan and Luma introduce slow-motion effects, and CogVideoX-1.5 and Open-Sora exhibit jittery, erratic movement. For a detailed analysis, see Appx. L.
@@ 4.3 Effect of Physics and Prompt Enhancement
Understanding real-world physics is essential for text-to-video models. Event-based prompts describe actions, while physics-enhanced prompts include natural consequences. A model’s ability to generate realistic videos based on these prompts reflects its real-world understanding. As proprietary models conceal their generation processes, we cannot verify prompt refinement. Instead, we analyze prompt types using open-source models, where prompt integrity can be manually ensured. In Tab. 4, we compare model performance across different prompt types to gauge the effects of physics-aware prompting and refinement. While refinement improves clarity, object detail, and aesthetics, it does not notably enhance physics accuracy. The persistent inconsistencies suggest that these refinements primarily address surface-level quality, leaving deeper physics limitations unresolved. On the other hand, physics-enhanced prompts often yield better results, reflecting the models’ strong prompt-following ability—even though they still struggle with true physical comprehension. Consequently, we propose a straightforward receipt: explicitly integrate physical phenomena into the prompt to nudge the model toward more realistic outcomes.
@@ 4.4 Anti-Physics Analysis
As shown in Tab. 3, all models show a performance decline from Fundamental to Composite to Anti-Physics categories, reflecting the challenge of generating complex physical phenomena. While they capture basic physics, simulating intricate interactions and “unphysical” scenes remains difficult, highlighting a gap in physics adherence. Our analysis found that failures often stem from models interpreting prompts in a more “reasonable” way rather than attempting and failing. For example, when prompted with “A glass sits on the table, and the wine in the glass has reversed gravity,” the model generates a still image, adhering to real-world physics instead of the prompt. This suggests that models prioritize mimicking training data over true physics understanding.
@@ 4.5 Failure to Handle Erratic Visual Changes
Certain physical events inherently result in erratic visual changes, such as collisions, breaking, or sudden state transitions. However, current video generation models struggle to accurately depict these scenarios. In the third example of Figure 10, a glass cup falls from a table, an event that should cause it to shatter upon impact. Instead, the model rationalizes the action, generating a video where the glass remains intact. Rather than simulating realistic disruptions in structure, motion, or object state, the models tend to default to continuous, smoothed animations, avoiding abrupt transformations that are critical for representing real-world physics. We created 20 paired prompts for each test, applied them to 12 video generation models (240 videos total), and only evaluate the visual change correctness through human evaluation. We also highlight Pika, the strongest model overall. Tab. 16 shows models struggle with abrupt visual changes — only 22.1% success overall vs. 42.1% for softened prompts; Pika 30% vs. 75%.When prompted with erratic visual changes, the model tends to simplify the task and generate the easier version.